OpenStudy (anonymous):

i am stuck on this problem. Square root of (x+deltaX)+ 2 - square root of (x-2)/ delta x

OpenStudy (anonymous):

\[\sqrt{(x+deltax)+2 } - \sqrt{(x+2) } \div deltax \]

OpenStudy (anonymous):

as delta x approaces 0

OpenStudy (anonymous):

Multiply top and bottom by sqrt(x+dx+2)+sqrt(x+2) Then try :P

OpenStudy (anonymous):

i did :D

OpenStudy (anonymous):

but i got lost like it was practically the same ting. i wasn't sure if anything can cancel out do i factor out something.

OpenStudy (anonymous):

Well you're supposed to get: \[\frac{1}{2\sqrt{x+2}}\]

OpenStudy (anonymous):

After you take the limit.

OpenStudy (anonymous):

i didnt get that.